how long will it take your php 16,000 to be double if invest at 8% simple interest
Answer by ikleyn(52794)

.

Write equation as you read the problem


    32000 = 16000 + 16000*0.08*t,


where t is the time (in years).  From the equation


    32000 - 16000 = 16000*0.08*t

        16000 = 16000*0.08*t

        1     =       0.08*t

        t     =       1/0.08 = 12.5.


ANSWER.  In 12.5 years.

Solved.